Please note: You are viewing the unstyled version of this website. Either your browser does not support CSS (cascading style sheets) or it has been disabled. Skip navigation.

June 24, 2014 - ESP Update

   Print Change text to small (default) Change text to medium Change text to large

All 3 ESPs have insufficient power now to run the hab assays. ESPjake was the last to go offline yesterday. ESP-1 (don) and ESP-2 (jake) have been reprogrammed and are currently acquiring CTD data through the end of June before mooring recovery in early July. ESP-3 (dennis) was also reprogrammed and was recording both CTD data and SUNA (nitrate) data, but we have now lost contact with ESPdennis, suggesting that there is no power left to transmit the CTD data. The others may follow suit...but have each completed their missions well.

We have attached the last plot of the ESP data for the 2014 bloom was very hard work, but also fun to provide the regional community with this service. Hope you found it helpful and interesting,

Without including a long list of people to thank...Please realize that MBARI stood with us closely here in the Anderson Lab at WHOI and really helped see each ESP mission through to its final completion...Thank-you Chris, Roman, and Brent for that excellent support.

Bruce Keafer

Last updated: July 1, 2019

whoi logo

Copyright ©2007 Woods Hole Oceanographic Institution, All Rights Reserved, Privacy Policy.
Problems or questions about the site, please contact